OpenStudy (anonymous):

find the linear approximation of the funtion f(x,y )= sqrt 20-x^2-7y^2 at (2,1). use your answer to aproximate f(1.95, 1.08)

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Linear_approximation a=2 b=1 then find partial derivatives wrt x and wrt y sub everything into formula

OpenStudy (anonymous):

f (2,1)+ sqrt 2x (x+2) + sqrt 14y ( y-1)

OpenStudy (dumbcow):

why is it sqrt(2x) ? is everything inside radical?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think is 1/2 ( 2x) ^-1/2

OpenStudy (anonymous):

1/ sqrt 2x

OpenStudy (dumbcow):

hmm \[d/dx \sqrt{20-x^{2}-7y^{2}} = \frac{-2x}{2\sqrt{20-x^{2}-7y^{2}}} = \frac{-x}{\sqrt{20-x^{2}-7y^{2}}}\]

OpenStudy (anonymous):

ok forget about that

OpenStudy (anonymous):

\[d/dy = -7y/\sqrt{20-x^2-14y^2}\]

OpenStudy (dumbcow):

right

OpenStudy (dumbcow):

oh should still be 7y^2 in radical

OpenStudy (dumbcow):

after plugging in (2,1) i get \[f(x,y) \approx 3-\frac{4}{3}(x-2)-\frac{7}{3}(y-1)\]

OpenStudy (anonymous):

@dumbcow after that i need to aproximate f(1.95, 1.08) doing the same thing that we did with (2,1)

OpenStudy (dumbcow):

no we already have the linear approximation plug in (1.95,1.08) for (x,y)

OpenStudy (anonymous):

@dumbcow ok